Weekly LIPAc highlights (14/June-18/June)


Plasma conditioning with H2 gas is regularly on-going while the vault is closed, with no issues.

Radio Frequency (RF) System and RFQ cavity

RF injection in the RFQ cavity and conditioning up to 10%DC is on-hold. It was identified that the low water mass flow rate in the HVPS (High Voltage Power Supplies) lines was caused by dirty diaphragms inside the flow regulator. The manifolds of the HVPS are being dismounted one by one and all the flow regulators are disassembled and carefully cleaned. The operation is almost completed: after reassembly on the HVPS a static pressure test and a circulation test are performed. All the HVPS cleaned so far confirmed that the nominal water mass flow rate is recovered. Actions are on-going in parallel to analyse the composition of the metallic residuals found to have clogged the flow regulators and to procure spare flow regulators as well as spare set of manifold + flow regulators. Cleaning by disassembly of flow regulators is a complex operation with high risk of breaking small parts: purchasing spares and even spare manifold is considered an important remedial action.

RF system is planned to be back operation during next week.

Medium Energy Beam Transport (MEBT)

The conditioning of the bunchers resumed compatibly with the concurrent intervention on the RF water cooling lines: the condition of buncher#1 progressed very quickly up to nominal cavity voltage (350 kV in CW for more than one hour).

High Energy Beam Transport (HEBT)

The debugging of the control system of the HEBT magnets PS has been completed with continuous support by CIEMAT with almost daily remote sessions.

The checkout of PPS signal related to the Lead Shutter was successfully completed.

Beam Dump

The water circulation including the Beam Dump cartridge in Mode 1 (main operational mode during beam operation) is stable including circulation in the BD cartiridge.

A massive operation is being carried out in the vault this Thursday and Friday with the reisntallation of the large radiation shield on the BD.

Beam Instrumentation

The stability of the accelerometers in the BD have been checked and confirmed in several remote sessions with CIEMAT.

Protection cover for RGBLM (Residual Gas Bunch Length Monitor) HV terminals have been installed in the vault and the conditioning for achieving the maximum voltage is on-going.

Control System

In view of the upcoming beam operation, the update of the required alarms to be implemented in the existing systems (BUZZER and BEAST) is on-going.


The regeneration of RFQ cryopumps (4 units out of  5) and HEBT cryopumps (pumps # 1, 2 and 3) was completed smoothly.

Installation of the Beam Dump shield in the vault

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.